OpenStudy (anonymous):

write the standard form of equation of the line decribe. through:(2,0), parallel to y=2/3x

OpenStudy (anonymous):

need you help

OpenStudy (aroub):

okay parallel lines have same slope , m=2/3x , (2,0) y=mx+b, 0=2/3(2) +b 0=4/3+b,3/3+4/3 =b , 7/3=b so, y=2/3x+7/3

OpenStudy (aroub):

the answer is y=2/3x+7/3

OpenStudy (anonymous):

ok thank you

OpenStudy (aroub):

ANYTIME! :D

OpenStudy (anonymous):

so you use formular y=mx+b?

OpenStudy (aroub):

YES

OpenStudy (aroub):

u can also use another formula

OpenStudy (aroub):

but i prefer to use y=mx+b

OpenStudy (anonymous):

what formular can be if you don't use last one?

OpenStudy (aroub):

ah sorry i really forgot how to use bec. i really dont use it but wait illl check in my book

OpenStudy (aroub):

if i find that book , im really sorry

OpenStudy (anonymous):

but its result is 2x-3y=4

OpenStudy (aroub):

okay my sis told me its y-1=m(x1-x1)

OpenStudy (aroub):

sorry y-y1=m(x-x1)

OpenStudy (aroub):

oh yes sry i just checked my answer its wrong its supposed to be y=3/2x-4/3

OpenStudy (anonymous):

what isvalue of m of this case

OpenStudy (aroub):

oh and 2x-3y=4 is the same as y=2/3x-4/3

OpenStudy (aroub):

m is the slope

OpenStudy (anonymous):

m =2/3 x?

OpenStudy (aroub):

2/3 without the x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if you use y-y1=m(x-x1, its equation is ?

OpenStudy (anonymous):

y1=0?

OpenStudy (anonymous):

x1=2?

OpenStudy (aroub):

same thing its just other equation to use exactly the same answer as the y=mx+b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you help me other one?

OpenStudy (aroub):

yes but its very slow in here wanna go to twiddla??

OpenStudy (aroub):

http://www.twiddla.com/570310

OpenStudy (anonymous):

throught:(-2,4), parallel to y=-3/2y +3

OpenStudy (anonymous):

ok thanks